OpenStudy (anonymous):

How do I make a rational equation with: If Beth takes one hour to bake a cake and it takes her daughter, Sally, an hour and 30 minutes to bake a cake, how long would it take for Beth and Sally to bake a cake together?

OpenStudy (anonymous):

lets call one hour and 30 minutes one and a half hours or better yet \(\frac{3}{2}\) hours

OpenStudy (anonymous):

then together they can bake the cake in \[\frac{1\times \frac{3}{2}}{1+\frac{3}{2}}\] hours

OpenStudy (anonymous):

we can take the steps if you like